The numbers of seats in the first 12 rows of a high school auditorium form an arithmetic sequence. The first row has nine seats. The second row has 11 seats.
a). Write a recursive formula to represent the sequence.
b). Write an explicit formula to represent the sequence.
c). How many seats are in the 12th row?

Answer by ramkikk66(644)

The numbers of seats in the first 12 rows of a high school auditorium form an arithmetic sequence. The first row has nine seats. The second row has 11 seats. 
a). Write a recursive formula to represent the sequence.
b). Write an explicit formula to represent the sequence.
c). How many seats are in the 12th row?

Ans:
a) Recursive formula for the number of seats in n'th row is: S(n) = S(n-1) + 2
b) Explicit formula is that of an arithmetic progression. If a is the first term, 
and d is the common difference, the nth term is given by a + (n - 1)*d
Here a is 9 and d is 2. Hence the number of seats in the n'th row will be
9 + 2*(n - 1)
c) Applying the above formula, number of seats in 12th row (i.e. n = 12)
is 9 + 2*(12 - 1) = 9 + 2*11 = 31 seats.

Hope this helps :)